Every element
x in mathbbZ has ____________ inverses with to respect to addition as a binary operation.
a.\\(-x\\)
b.\\(0\\)
c.\\(frac{1}{x}\\)
d.\\(x\\)
1
Expert's answer
2020-05-20T18:55:52-0400

First we will find the identity of "Z" with respect to addition.

There exist "0 \in Z" called identity of "Z" such that "x+0=x". So, "e=0" is identity of "Z" with respect to addition.

Now, let "x" be any element of "Z", then "y" be the inverse of "x" if "x+y=y+x = 0"

"\implies y = -x" and since "x\in z \implies y=-x \in Z".

So, "-x" is the inverse of "x".

Correct option is (a) "-x" .